OpenStudy (anonymous):

The functions f(x) and g(x) are described using the following equation and table: f(x) = -3(1.02)x x g(x) -1 -5 0 -3 1 -1 2 1 Which statement best compares the y-intercepts of f(x) and g(x)?

OpenStudy (anonymous):

@DanJS lemme rewrite that.

OpenStudy (anonymous):

x g(x) -1 -5 0 -3 1 -1 2 1

OpenStudy (anonymous):

The y-intercept of f(x) is equal to the y-intercept of g(x). The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). The y-intercept of g(x) is equal to 2 plus the y-intercept of f(x).

OpenStudy (danjs):

k

OpenStudy (danjs):

So . f(x) = -3(1.02)^x

OpenStudy (danjs):

and g(x) is the table of values?

OpenStudy (danjs):

The Y-intercept is when X=0 f(0) = -3(1.02)^0 = -3(1) = -3 Y-intercept of f(x) is y = -3

OpenStudy (danjs):

From the table for g(x) g(0) = -3

OpenStudy (danjs):

The y intercepts are the same for both g(x) and f(x), y=-3

OpenStudy (anonymous):

Okay :)